OpenStudy (anonymous):

what is the value of x2-x+3?

OpenStudy (anonymous):

for which value of x?

OpenStudy (anonymous):

yea? which value

OpenStudy (anonymous):

hold on i typed the question wrong

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{ 2x-3 }{ x }=x+1\Rightarrow x^2+x=2x-3 \Rightarrow x^2-x+3=0\] )

OpenStudy (anonymous):

so its 0

OpenStudy (anonymous):

you tell me...

OpenStudy (anonymous):

can u walk me through the steps

OpenStudy (anonymous):

sure... first multiply both sides by x to get rid of the fraction\[\left( \frac{ 2x-3 }{ x } \right)\cdot x=\left( x+1 \right)\cdot x \Rightarrow 2x-3=x^2+x\] then subtract (2x-3) from both sides.\[\begin{matrix}2x-3 =x^2+x\\ -2x+3 \,\,\,-2x+3 \\\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,0=x^2-x+3\end{matrix}\] you end up with 0.

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

yea it5 makes perfect sense

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

one thing you should be aware of... x is originally in the denominator. we can't divide by 0 so we must insist that \(x\ne 0\). if at some point we solve for x and find that 0 is a solution to the analogous equation, we must exclude it from the solution set.

OpenStudy (anonymous):

however, for this particular equation, x=0 is not a solution.

OpenStudy (anonymous):

yea i know but thanks alot

OpenStudy (anonymous):

you're welcome